Assume that verbal scores on the SAT exam have a normal distribution with a population mean of 500 points and a population standard deviation of 100. 8. What is the probability that a randomly chosen makes at least 510 but no more than 620? a. 0.2145 b. 0.2422 c. 0.2850 d. 0.3451 e. 0.4247
Answer #1

Solution :

=> Answer : Option d. 0.3451

Given that mean μ = 500 , standard deviation σ = 100

=> P(at least 510 but no more than 620) = P(x >= 510 , x <= 620)

= P(510 <= x <= 620)

= P(509.5 < x < 620.5)

= P((509.5 - 500)/100 < (x - μ)/σ < (620.5 - 500)/100)

= P(0.095 < Z < 1.205)

= 0.3451